Solve the system of equation:
3x+2y=11
2x-2y=4
Solve the system:
2x-y+2z=11
3x-2y-z=12
x-3y-z=1

Answer by mangopeeler07(462)

3x+2y=11
2x-2y=4

Add the equations together to eliminate the y's.
5x=15.

Divide both sides by 5.
x=3.

Plug 3 in for x in either of the original equations.

2(3)-2y=4
6-2y=4
y=1.

So x=3 and y=1.